# score-based generative model SNR signal-to-noise ratio

### v2

@@
Perceptual Objectve Listening Quality Analysis
@@
Julius Richter , , Simon Welker , , Jean-Marie Lemercier , , Bunlong Lay , Timo Gerkmann This work has been funded by the German Research Foundation (DFG) in the transregio project Crossmodal Learning (TRR 169), DASHH (Data Science in Hamburg - HELMHOLTZ Graduate School for the Structure of Matter) with the Grant-No. HIDSS-0002, and the Federal Ministry for Economic Affairs and Climate Action, project 01MK20012S, AP380. We would like to thank J. Berger and Rohde&Schwarz SwissQual AG for their support with POLQA.Simon Welker is with the Signal Processing Group, Department of Informatics, Universität Hamburg, 22527 Hamburg Germany, and with the Center for Free-Electron Laser Science, DESY, 22607 Hamburg, Germany (e-mail: simon.welker@uni-hamburg.de). The other authors are all with the Signal Processing Group, Department of Informatics, Universität Hamburg, 22527 Hamburg Germany (e-mail: {julius.richter; jeanmarie.lemercier; bunlong.lay; timo.gerkmann}@uni-hamburg.de). $0$-0002-7870-4839 $0$-0002-6349-8462 $0$-0002-8704-7658 $0$-0002-0847-7896 $0$-0002-8678-4699
@@ Abstract
Index Terms:
speech enhancement, dereverberation, diffusion models, score-based generative models, score matching.
{tikzpicture}
[…]
@@ I Introduction
Nonetheless, note that diffusion-based speech enhancement methods, unlike the VAE-based method described above, are not counted as unsupervised methods, since labeled data (i.e. clean and noisy speech pairs) are used for training. However, the learning objective remains generative in nature which is to learn a prior for clean speech per se rather than a direct mapping from noisy to clean speech. In fact, supervision is only exploited to learn the conditional generation of clean speech when noisy speech is given. Thus, current diffusion-based models for speech enhancement, such as [23, 21, 22], can be considered conditional generative models trained in a supervised manner.
@@ II-A Data representation
Thus, we operate on complex spectrograms that are elements of $\mathbb{C}^{T\times F}$, where $T$ denotes the number of time frames dependent on the audio length, and $F$ represents the number of frequency bins. To compensate for the typically heavy-tailed distribution of STFT speech amplitudes [33], we apply an amplitude transformation
@@
to all complex STFT coefficients $c$, where ($\angle$($\cdot$)∠ ( ⋅ ) represents the angle of a complex number, $\alpha\in(0,1]$ is a compression exponent which brings out frequency components with lower energy (e.g. fricative sounds of unvoiced speech) [34], and $\beta\in\mathbb{R}_{+}$ is a simple scaling factor to normalize amplitudes roughly to within $[0,1]$. Such a compression has been argued to be perceptually more meaningful in speech enhancement [35, 36], and the transformation ensures that the neural network operates on consistently scaled inputs with respect to the Gaussian diffusion noise [25].
@@ II-B Stochastic Process
The tasks at hand, speech enhancement and dereverberation, can be considered as conditional generation tasks: Given the corrupted noisy/reverberant speech, generate clean speech by using a conditional generative model. Most previously published diffusion-based generative models are adapted to such conditional tasks either through explicit conditining channels added to the DNN [37, 38], or through combining an unconditionally trained score model with a separate model (such as a classifier) that provides conditioning in the form of a gradient [39, 28]. With our method, we explore a third possibility, which is to incorporate the particular task directly into the forward and reverse processes of a diffusion-based generative model.
@@
where $\mathbf{x}_{t}$ is the current process state, $t\in[0,T]$ a continuous time-step variable describing the progress of the process (not to be confused with the time index of any signal in the time or T-F domain), $\mathbf{y}$ the noisy or reverberant speech, and $\mathbf{w}$ denotes a standard Wiener process. The vector-valued function $\mathbf{f}(\mathbf{x}_{t},\mathbf{y})$ is referred to as the drift coefficient, while $g(t)$ is called the diffusion coefficient and controls the amount of Gaussian white noise injected at each time-step. Note that different to Song et al. [28], our drift term is now a function of $\mathbf{y}$, by which we tailor the proposed SDE to reconstruction tasks. The process is defined for each T-F bin independently. Thus, the variables in bold are assumed to be vectors in $\mathbb{C}^{d}$ with $d=TF$ containing the coefficients of a flattened complex spectrogram.
@@
$\mathrm{d}{\mathbf{x}_{t}}=\left[-\mathbf{f}(\mathbf{x}_{t},\mathbf{y})+g(t)^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})\right]\mathrm{d}{t}+g(t)\mathrm{d}{\bar{\mathbf{w}}}\,,$ (5)
@@
$\mathrm{d}{\mathbf{x}_{t}}=\left[-\mathbf{f}(\mathbf{x}_{t},\mathbf{y})+g(t)^{2}\mathbf{s}_{\theta}(\mathbf{x}_{t},\mathbf{y},t)\right]\mathrm{d}{t}+g(t)\mathrm{d}{\bar{\mathbf{w}}}\,,$ (6)
@@ II-C Training objective
$\nabla_{\mathbf{x}_{t}}$ $\log p_{0t}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{y})=\nabla_{\mathbf{x}_{t}}\log\left[|2\pi\sigma\mathbf{I}|^{-\frac{1}{2}}\mathrm{e}^{-\frac{\|\mathbf{x}_{t}-\boldsymbol{\mu}\|^{2}_{2}}{2\sigma^{2}}}\right]$ (11)
@@ III Numerical SDE Solvers
$\mathrm{d}{\mathbf{x}_{t}}=\left[-\mathbf{f}(\mathbf{x}_{t},\mathbf{y})+g(t)^{2}\mathbf{s}_{\theta}(\mathbf{x}_{t},\mathbf{y},t)\right]\mathrm{d}{t}\,,$ (16)
@@ V-D Hyperparameters and training configuration
[…]
@@ V-E Sampler settings
[…]
@@
ODE atol=$10^{-1}$, rtol=$10^{-1}$ 14 0.46 2.78 12.83
ODE atol=$10^{-6}$, rtol=$10^{-3}$ 49 1.55 2.71 12.76
(a) Varying $N$, fix $r=0.33$
@@ VI-A Speech Enhancement
[…]
@@
Finally, Tab. III lists the results for the standardized VB-DMD dataset. This has the advantage that one can take values from other methods and copy them from the corresponding papers for a quick algorithmic comparison. It can be seen that SGMSE+ outperforms all other generative baselines, further narrowing the performance gap with discriminative methods that currently lead the benchmark based on PESQ [^3] https://paperswithcode.com/sota/speech-enhancement-on-demand, including recent approaches such as [79] and [80]. It should however be noted that PESQ formally requires a minimum file length of 3.2 sec according to P.862.3 [81], which is not the case for most files in VB-DMD [53].
TABLE III: Speech enhancement results obtained for VB-DMD. Models marked with an asterisk ($*{}^{*}$) are additional baselines with values taken from the corresponding papers.

### v3

@@
Perceptual Objective Listening Quality Analysis
@@
Julius Richter , , Simon Welker , , Jean-Marie Lemercier , , Bunlong Lay , Timo Gerkmann  This work has been funded by the German Research Foundation (DFG) in the transregio project Crossmodal Learning (TRR 169), DASHH (Data Science in Hamburg - HELMHOLTZ Graduate School for the Structure of Matter) with the Grant-No. HIDSS-0002, and the Federal Ministry for Economic Affairs and Climate Action, project 01MK20012S, AP380. We would like to thank J. Berger and Rohde&Schwarz SwissQual AG for their support with POLQA.Simon Welker is with the Signal Processing Group, Department of Informatics, Universität Hamburg, 22527 Hamburg Germany, and with the Center for Free-Electron Laser Science, DESY, 22607 Hamburg, Germany (e-mail: simon.welker@uni-hamburg.de). The other authors are all with the Signal Processing Group, Department of Informatics, Universität Hamburg, 22527 Hamburg Germany (e-mail: {julius.richter; jeanmarie.lemercier; bunlong.lay; timo.gerkmann}@uni-hamburg.de).
@@ I Introduction
Nonetheless, note that diffusion-based speech enhancement methods, unlike the VAE-based method described above, are not counted as unsupervised methods, since labeled data (i.e. clean and noisy speech pairs) are used for training. However, the learning objective remains generative in nature which is to learn a prior for clean speech per se rather than a direct mapping from noisy to clean speech. In fact, supervision is only exploited to learn the conditional generation of clean speech when noisy speech is given. Thus, current diffusion-based models for speech enhancement, such as [23, 21, 22], can be considered as conditional generative models trained in a supervised manner.
@@ II-A Data representation
Thus, we operate on complex spectrograms that are elements of $\mathbb{C}^{K\times F}$, where $K$ denotes the number of time frames dependent on the audio length, and $F$ represents the number of frequency bins. To compensate for the typically heavy-tailed distribution of STFT speech amplitudes [33], we apply an amplitude transformation
@@
to all complex STFT coefficients $c$, where $\angle(\cdot)$ represents the angle of a complex number, $\alpha\in(0,1]$ is a compression exponent which brings out frequency components with lower energy (e.g. fricative sounds of unvoiced speech) [34], and $\beta\in\mathbb{R}_{+}$ is a simple scaling factor to normalize amplitudes roughly to within $[0,1]$. Such a compression has been argued to be perceptually more meaningful in speech enhancement [35, 36], and the transformation ensures that the neural network operates on consistently scaled inputs with respect to the Gaussian diffusion noise [25].
@@ II-B Stochastic Process
The tasks at hand, speech enhancement and dereverberation, can be considered as conditional generation tasks: Given the corrupted noisy/reverberant speech, generate clean speech by using a conditional generative model. Most previously published diffusion-based generative models are adapted to such conditional tasks either through explicit conditioning channels added to the DNN [37, 38], or through combining an unconditionally trained score model with a separate model (such as a classifier) that provides conditioning in the form of a gradient [39, 28]. With our method, we explore a third possibility, which is to incorporate the particular task directly into the forward and reverse processes of a diffusion-based generative model.
@@
where $\mathbf{x}_{t}$ is the current process state, $t\in[0,T]$ a continuous time-step variable describing the progress of the process (not to be confused with the time index of any signal in the time or T-F domain), $\mathbf{y}$ the noisy or reverberant speech, and $\mathbf{w}$ denotes a standard Wiener process. The vector-valued function $\mathbf{f}(\mathbf{x}_{t},\mathbf{y})$ is referred to as the drift coefficient, while $g(t)$ is called the diffusion coefficient and controls the amount of Gaussian white noise injected at each time-step. Note that different to Song et al. [28], our drift term is now a function of $\mathbf{y}$, by which we tailor the proposed SDE to reconstruction tasks. The process is defined for each T-F bin independently. Thus, the variables in bold are assumed to be vectors in $\mathbb{C}^{d}$ with $d=KF$ containing the coefficients of a flattened complex spectrogram.
@@
$\hskip-1.90002pt\mathrm{d}{\mathbf{x}_{t}}=\left[\mathbf{f}(\mathbf{x}_{t},\mathbf{y})-g(t)^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})\right]\mathrm{d}{t}+g(t)\mathrm{d}{\bar{\mathbf{w}}}\,$, (5)
@@
$\mathrm{d}{\mathbf{x}_{t}}=\left[\mathbf{f}(\mathbf{x}_{t},\mathbf{y})-g(t)^{2}\mathbf{s}_{\theta}(\mathbf{x}_{t},\mathbf{y},t)\right]\mathrm{d}{t}+g(t)\mathrm{d}{\bar{\mathbf{w}}}\,$, (6)
@@ II-C Training objective
$\hskip-1.49994pt\nabla_{\mathbf{x}_{t}}$ $\log p_{0t}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{y})=\nabla_{\mathbf{x}_{t}}\log\left[|2\pi\sigma\mathbf{I}|^{-\frac{1}{2}}\mathrm{e}^{-\frac{\|\mathbf{x}_{t}-\boldsymbol{\mu}\|^{2}_{2}}{2\sigma^{2}}}\right]$ (11)
@@ III Numerical SDE Solvers
$\mathrm{d}{\mathbf{x}_{t}}=\left[\mathbf{f}(\mathbf{x}_{t},\mathbf{y})-\frac{1}{2}g(t)^{2}\mathbf{s}_{\theta}(\mathbf{x}_{t},\mathbf{y},t)\right]\mathrm{d}{t}\,$, (16)
@@ V-D Hyperparameters and training configuration
We convert each audio input with sampling rate $16\,\text{kHz}$ into a complex-valued STFT representation using a window size of 510, resulting in $F=256$, a hop length of 128 (i.e. approximately 75% overlap), and a periodic Hann window. To process multiple examples for batch training, the length of each spectrogram is trimmed to $K=256$ STFT time frames, with start and end times selected randomly at each training step. For the spectrogram transformation in Eq. (1), we have chosen $\alpha=0.5$ and $\beta=0.15$ empirically.
@@ V-E Sampler settings
TABLE I: Results for different sampler configurations tested on VB-DMD with the average number of function evaluations (NFE) and the respective average real-time factor (RTF)2.
@@
[…]
(a) Varying NN, fix $r=0.33$
@@ VI-A Speech Enhancement
In Tab. II, we report the speech enhancement results on the WSJ0-CHiME3 test set for the matched and mismatched condition, i.e. when the training set was also WSJ0-CHiME3 or when the training set was VB-DMD. We compare our proposed method, which we call SGMSE+, with selected baseline methods and sort the results by the type of algorithm, which is either generative or discriminative. Considering the matched condition in the upper half of Tab. II, we see that SGMSE+ outperforms all other generative methods in all metrics. Note that STCN and RVAE are both unsupervised speech enhancement methods, i.e. they are trained on clean speech only (WSJ0 or VB). RVAE shows competitive results for SI-SAR, however, its VEM optimization algorithm is very time-consuming due to the fine-tuning of the encoder at test time, resulting in a RTF of >>10000. This is significant in contrast to STCN with a RTF of 0.64 and SGMSE+ with a RTF of 1.772. Although both VAE-based methods model temporal dependencies, they are limited in their ability to produce high-quality speech, likely due to the dimensionality reduction of the latent variable and the encoder’s sensitivity to noisy input, which causes the latent variable to be incorrectly initialized [15].
@@
Finally, Tab. III lists the results for the standardized VB-DMD dataset. This has the advantage that one can take values from other methods and copy them from the corresponding papers for a quick algorithmic comparison. It can be seen that SGMSE+ outperforms all other generative baselines, further narrowing the performance gap with discriminative methods that currently lead the benchmark based on PESQ, including recent approaches such as [79] and [80]. It should however be noted that PESQ formally requires a minimum file length of 3.2 sec according to P.862.3 [81], which is not the case for most files in VB-DMD [53].
TABLE III: Speech enhancement results obtained for VB-DMD. Models marked with an asterisk () are additional baselines with values taken from the corresponding papers.